# Pauli Check Sandwiching for Quantum Characterization and Error Mitigation during Runtime

### v2

@@ Abstract
This work presents a novel quantum system characterization and error mitigation framework that applies Pauli check sandwiching (PCS). We motivate our work with prior art in software optimizations for quantum programs like noise-adaptive mapping and multi-programming, and we introduce the concept of PCS while emphasizing design considerations for its practical use. We show that by carefully embedding Pauli checks within a target application (i.e. a quantum circuit), we can learn quantum system noise profiles. Further, PCS combined with multi-programming unlocks non-trivial fidelity improvements in quantum program outcomes.
Index Terms:
quantum computing, quantum error mitigation, quantum device characterization, quantum software.
@@ I Introduction and Motivation
The development of a practical quantum computer (QC) is highly anticipated due to its potential to provide significant speedups for certain classes of high-complexity problems that are intractable today, such as those in chemistry and optimization. While quantum hardware and software stacks have improved significantly over the past decade, existing quantum system noise still varies drastically cross-chip. An example is illustrated in Fig. 1. In many QCs existing today, the magnitude of average error hovers above error correction thresholds. Additionally, the demand of QCs, largely hosted in the quantum cloud [15], greatly outnumbers supply, inflating costs to access hardware.
@@
Much work has been dedicated to improving both quantum program success and machine utilization during runtime in an effort to ease bottlenecks that hinder scaling. For example, noise-adaptive mapping is a state-of-art quantum circuit optimization technique used to place and route quantum programs on quantum hardware [12]. Noise-aware schemes, however, depend on accurate calibration data from the QC provider or from benchmarking. Even in the case that a snapshot of machine properties were available, QC noise is observed to fluctuate over time in ways that are challenging to capture with analytical models[2]. If used for hardware allocation, sub-optimal mappings result.
Figure 1: Example QPU with 20 regions (10 qubits per region) of unique 1q error rates, $p\in[0.0005,0.01]$, and 2q errors, $2p$. PCS-based characterization with multi-programmed circuits learns the QPU’s previously unknown noise landscape.
@@
In our work, we seek alternative methods to determine the noise profile of a QC when its properties are unknown. It is desirable to develop characterization techniques with minimal overheads in terms of quantum processor unit (QPU) runtime. We are inspired by prior art in multi-programming [1] and ensemble distributions [16] to explore how parallelization during runtime can be employed for 1) learning quantum noise or hardware characterization, 2) fidelity improvements in program outcomes, and 3) reductions in runtime latency. This work presents a novel framework that combines the power of multi-programming with Pauli check sandwiching (PCS) for QC characterization and error mitigation. Multi-programming is a technique for parallelized quantum circuit execution while PCS is an error mitigation strategy that utilizes post-selection to reduce errors in quantum circuit outcomes. Our techniques intelligently embed PCS into quantum applications. These programs are then simultaneously executed across a QPU for noise profile characterization. Resulting characterization data adaptively mitigates noisy outcomes within individual thread outcomes. Knowledge of the QPU’s best regions creates a weighted ensemble result with non-trivial gain, found to be up to a $25\%$ fidelity improvement in our reported results.
We speculate that the methods presented here will be advantageous in the not-too-distant future – as larger quantum processors are deployed on the quantum cloud, quantum resource providers may not provide fine-grained particulars about QC operational characteristics and error rates as we are familiar with today. Additionally, our techniques are incredibly relevant to dynamic defects that are observed to spike local infidelity rates on quantum platforms during computation. An example of a dynamic defect would be a cosmic ray strike multi-bit burst errors [10]. We also note that various other factors can cause qubit properties and errors to drift.
II Prior Art in Quantum Program Optimization
QC hardware is noisy during use and expensive to access. As a result, much work has been dedicated to improving both quantum program success and machine utilization during runtime. This section will discuss QC noise-adaptive mapping and multi-programming, two quantum program optimization techniques that push the frontier of existing hardware while presenting new questions that motivate this work.
This section does not provide an exhaustive introduction to quantum computation. We recommend that the reader refers to Ref. [13] for a more complete introduction to topics in quantum information processing and Ref. [4] for an overview of the fundamentals in quantum computing systems.
II-A Noise-adaptive Mapping
There are a variety of different qubit technologies that exist today, such as superconducting circuits, trapped ions, and neutral atoms, among others. Each of these platforms have their respective benefits and costs, and they all demonstrate a wide range of performance characteristics in terms of fidelity, coherence, and gate times. Even QCs comprised of the same qubit technology tend to demonstrate performance differences. Due to a combination of imperfect fabrication procedures and intrinsic environmental conditions, each quantum system is characterized by a unique noise landscape during computation. This presents the challenge of determining how to best allocate logical qubits to physical qubits during runtime.
Noise-adaptive mapping is a state-of-art mapping technique used to place and route quantum programs on quantum hardware [17, 12]. Noise-aware schemes, however, often depend on the availability of calibration data from the QC service provider. Further, these schemes also assume that the calibration data provided by the quantum hardware vendor is complete and correct. If trusted machine data is unavailable, protocols such as SABRE [8] will be left without (or with misinformed) guidance for logical to physical qubit mapping, resulting in poor quantum program performance. Even in the case that machine characterization data is complete and derives from a reliable source, it is known that machine noise and performance characteristics have a tendency to fluctuate over time in ways that are challenging to capture with analytical models[2]. If stale calibration data is used in this case, sub-optimal mappings result.
In our work, we explore alternative methods to determine the noise profile of quantum hardware due to the challenges associated with depending on provider-supplied data. It is desirable to develop techniques with minimal overheads associated with gathering characterization data. In a work with similar motivations as our own, Ref. [14] attempts to characterize the noise that an application will experience on a target QPU through the use of easy-to-simulate Clifford-circuits that approximate the structure of the circuit of interest. However, this Clifford-based technique has drawbacks. First, additional QC runtime is required to gather characterization data from the Clifford circuits. Second, the actual payload circuits must be run soon after characterization data is retrieved to avoid sub-optimal results from machine property drift. These two drawbacks inspire us to develop techniques that allow characterization routines to be interleaved into the actual circuits of interest to avoid additional QC benchmarking and its necessary runtime.
II-B Multi-programming
Some QCs on today’s quantum cloud contain more than 100 qubits, but low qubit coherence times and high operator error prohibits the use of all of these qubits collectively for computation. As a result, QC applications are typically much smaller than the qubit capacity of the machine, leaving many QC qubits unused during runtime. Recent work has explored how to take advantage of these unused qubit resources to improve quantum program outcomes.
Ideally, compute resources are scheduled for maximum utilization. As the work in Ref. [19] highlights, parallelized quantum circuit execution, as compared to a sequential approach, offers the benefit of reduced runtime latency because a quantum circuit executed $n$ times simultaneously can reach a threshold number of observables in a distribution in $1/n$ time. However, work in Ref. [1] highlights that multi-programming quantum hardware, especially machines existing today, is a non-trivial task as improper scheduling procedures can cause undesirable crosstalk between two different workloads. The observation of crosstalk is especially likely whenever two quantum programs of different duration are executed adjacently – the measurement operations of the shorter circuit could impact the longer circuit that still has quantum circuit operations in-progress. If implemented with care, however, multi-programming can increase quantum system throughput without having a detrimental impact on program outcomes.
We are interested in exploring the potential that multi-programming has for improving quantum program performance. The work in [16] demonstrates how ensembling techniques employed to create a distribution from many circuit mappings can help smooth errors, improving fidelity. Creating a cumulative distribution from many instances of the same application implemented in parallel could have a similar effect that boosts performance.
III Pauli Check Sandwiching
We rely on a technique called ‘Pauli Check Sandwiching’ (PCS) [5, 3] to detect and mitigate errors. Prior work in PCS has shown its viability for mitigating gate and measurement errors in quantum circuits [9] and for fault injection [11] that studies error in quantum circuits. In our work, we also find that PCS is a powerful tool for guiding QC characterization, and we make the novel contribution of combining PCS with multi-programming.
@@ III Pauli Check Sandwiching
As seen in Fig. 2, PCS surrounds a payload circuit, $U$, with controlled Pauli operator checks. Errors on $U$ can be detected on an ancilla through phase kickback. It is important that the relationship
$R_{1}UL_{1}=U$ (1)
@@
Figure 3: Benchmarks used for analysis of methods included (a) an eight-qubit GHZ / sensing circuit and (b) the three-qubit Toffoli operation. PCS checks are circled in red. The portion of each circuit protected by the PCS checks, $U$, lie in between the circled checks. Two additional ancilla qubits were needed for the two single-qubit PCS checks in both (a) and (b).
holds so the introduced checks do not disturb the original payload circuit semantics. As a note, the operations contained in $L_{1}$ and $R_{1}$ do not have to comprise the same Pauli gates as long as Eqn. 1 is satisfied.
Recognizing that all errors can be decomposed into Pauli operators is a key concept in understanding PCS. By cleverly constructing Pauli sandwich checks dependent on $U$, anti-commutativity relations between the errors and checks can reveal errors in the ancilla qubits. Measuring 1 in any ancillas indicates that phase kickback occurred, meaning anti-commutation between the Paulis, and thus, a present error. As part of the PCS protocol, we discard this error-corrupted shot, leading to an increase in final distribution fidelity of our circuit of interest, $U$.
Practical application of PCS requires consideration of tradeoffs. First, determining the Pauli unitary used in the check increases in complexity with the circuit. Second, Pauli check logic must be decomposed into the supported gate set of the QC, and as a result, large multi-qubit gates will end up being high-depth. Third, the overhead of adding PCS into the circuit must not outweigh the corrective benefit in terms of gate error. Fourth, while increasing the number of check pairs improves error detection, it also introduces additional gate noise to the circuit and exponentially increases the required sampling size. We note that the exploration of an application’s ‘check limit’ is included in Ref. [7]. All of these limitations must be considered for practical implementation of PCS within quantum computing workflows.
IV QC Characterization and Error Mitigation with PCS
The previous section describes how the check data gathered during PCS by measuring a Pauli check’s ancilla qubit can inform whether or not an error occurred during the execution of a sandwiched circuit, $U$. Over many circuit samples, the amount of times that an ancilla is measured to be 1 can provide insight to the extent of a quantum system’s noise. Consider the case where two identical circuits embedded with Pauli checks are executed in parallel on different regions of a QPU. If the check data is compared between two circuits, we can infer which execution contained more noise since a higher count of checks measuring as 1 indicates more error. Thus, we propose that by combining multi-programming with the PCS protocol, we can 1) estimate the noise profile of a target QC and 2) increase circuit fidelity by removing the measured outcomes that contain errors.
Many of today’s QPUs feature more qubits than are required for a target application. In this situation, numerous mapping permutations that assign logical to physical qubits are possible, and this observation introduces the question of how to make the best use of a quantum device. On one hand, one could prioritize using the best regions on-chip, making circuit execution serialized in order to strategically target the high-performing regions. However, this approach requires a priori knowledge of the properties of the QPU. In addition, serialized execution necessitates longer periods of dedicated hardware access. On the other hand, multi-programming to generate an ensemble distribution amplifies throughput via parallelization, potentially enabling lower-latency runtime, but it comes at the cost of including results from less reliable qubits if the error profile varies on the QPU.
In our work, we assume that the number of qubits on chip is larger than the number of logical qubits required for an application. Further, we assume no prior knowledge of the noise profile on-chip, and we attempt to maximize hardware utilization. We inject PCS into our target application, but only single-qubit checks on edge qubits are employed to minimize PCS gate count overhead. To maximally utilize hardware via multi-programming, a QC with $n$ physical qubits holds
$threads=\lfloor n/(q_{algorithm}+q_{ancilla})\rfloor$ (2)
instances of a circuit running in parallel. In this equation, $q$ indicates qubits, both algorithm and ancilla, that are in the logical quantum circuit.
Assuming the same shot budget in each each local thread, values of the checks (i.e. PCS ancilla measurements of 1) are used during post processing to filter error-corrupted outcomes from each local thread’s measurement distribution. This process creates a vector of error-mitigated counts for each thread, $c_{i}$. Circuits that encounter more noise will discard a greater count of shots. Discarded shot count for each local thread is represented with the scalar $r_{i}$. The percentage of discarded measurement outcomes for each thread,
$d_{i}=\frac{r_{i}}{shots_{i}},$ (3)
gives insight into the underlying noise profile of the QC when different regions are compared. Each local $d_{i}$ is then used as a weight that scales $c_{i}$ to create $s_{i}$:
$s_{i}=c_{i}*\frac{min(d)}{d_{i}}.$ (4)
In Eqn. 4, $min(d)$ represents the minimum percentage of discarded shots from all the parallel threads. A cumulative distribution, $S$, is created from the sum of the $s_{i}$ vectors from the PCS-protected local threads,
$S=\sum s_{i}.$ (5)
V Experiments and Analysis
In this section, we evaluate our methods for applying PCS toward quantum characterization and error mitigation. During our experiments, we targeted two applications. Our first application was an eight-qubit GHZ mirror circuit that could be used for evaluating quality of multi-qubit entanglement as well as quantum sensing applications [18]. Our second application was a three-qubit Toffoli gate since this operation, especially its generalized form, frequently appears in quantum computing applications. The circuit structures for our benchmarks appears in Fig. 3. In both circuits, we protect only the edge qubits to introduce less overhead. We inject PCS that use the same Pauli operation (i.e. $L_{1},R_{1}=X$ or $L_{1},R_{1}=Z$) to simplify the demonstration of our techniques. However, we note that this constraint can be relaxed as that Pauli checks do not need to be the same unitary for PCS to be effective. For these benchmarks, it is probable that more optimal checks exist that provide more robust error mitigation. We note that discovering optimal checks has poor scaling as $L_{1}$ and $R_{1}$ span more qubits.
Our experimental results were generated using noisy simulation implemented with Qiskit Aer [6]. The test bench used during evaluation was an expansion of the model pictured in Fig. 1 to emulate a QC with 60 regions where single-qubit depolarizing noise $p\in[0.0005,0.03]$ and two-qubit gates $2p$. One instance of each application was allocated per region so that 60 threads of each benchmark executed in parallel. The circuits pictured in Fig. 3 were transpiled to use the basis gate set that included $CX$, $X$, $SX$, and $RZ$.
Fig. 4 shows our results, detailing a multiple percentage point increase in fidelity when multi-programmed applications are protected with PCS and a cumulative distribution, calculated with Eqn. 5, is created from local results weighted by check data. The baseline used for comparison in Fig. 4 is an ensemble distribution formed by multi-programmed benchmarks that do not contain PCS. Even with only the edge qubits protected by Pauli checks, a measurable benefit results from our approach that combines PCS and multi-programming.
Fig. 5 documents the relationship between discarded shots and error rate, illustrating PCS’s potential for QC error characterization. Further, using this relationship to predict the noise profile of a QC chip, as shown in Fig. 6 and Fig. 7, can be extremely accurate. By using the percentage of discarded shots in each region, we can estimate relative underlying error rates for those regions. This gives us further indication that PCS can be used in conjunction with multi-programming to scope out the noise profile of a QC, pushing towards higher fidelity results for quantum applications.
@@ V Experiments and Analysis
Figure 4: Differences between fidelity measurements of a circuit with no PCS (base) vs. PCS protection. 10,000 shots were used for both circuits, and fidelity improvement was determined to be $25\%$ (left) and $18.75\%$ (right).
@@
Figure 5: Shots removed / total shots as error rates increase. 10,000 shots on a GHZ circuit, pictured in Fig. 3(a), were executed for every error rate $p\in[0.0005,0.03]$ for one-qubit gates and two-qubit gates $2p$ (60 total noise models).
Figure 6: Heatmap of an example QC chip noise profile, where each square is a 10 qubit region of the chip characterized by depolarizing noise, $p$.
Figure 7: Heatmap of plotting the percentage of shots removed after running 10,000 shots on PCS-protected GHZ circuit on each region. The underlying QC chip is the same as Fig. 6.
VI Conclusion and Future Work
Combining PCS with multi-programming creates a framework that boosts quantum program performance. We find that PCS not only mitigates error in individual circuit distributions – check data can also be used to intelligently combine the results of multiple threads cross-QPU, demonstrating measurable noise reductions in final outcomes. But even more powerfully, PCS has additional promise to guide efficient qubit mapping when used in conjunction with multi-programming. When lacking reliable data on the noise profile of a QC, it is important that we find methods to best uncover it, guiding hardware utilization, especially if a shot budget must be considered. We theorize that running PCS circuits on different sections of a quantum chip in parallel could enable real-time qubit mapping with check data. This “noise-unaware” mapping is in development.
VII Acknowledgments

### v3

@@ Abstract
This work presents a novel quantum system characterization and error mitigation framework that applies Pauli check sandwiching (PCS). We motivate our work with prior art in software optimizations for quantum programs like noise-adaptive mapping and multi-programming, and we introduce the concept of PCS while emphasizing design considerations for its practical use. We show that by carefully embedding Pauli checks within a target application (i.e. a quantum circuit), we can learn quantum system noise profiles. Further, PCS combined with multi-programming unlocks non-trivial fidelity improvements.
@@ I Introduction and Motivation
While the quantum compute stack has improved, existing quantum system noise still varies cross-chip (Fig. 1) and hovers above error correction thresholds. Much work has been dedicated to improving quantum program success. For example, noise-adaptive mapping is a state-of-art quantum circuit optimization technique used to place and route quantum programs on quantum hardware  [5]. Noise-aware schemes, however, depend on accurate calibration data from the quantum computer (QC) provider or from benchmarking. Even in the case that a snapshot of machine properties were available, QC noise is observed to fluctuate over time in ways that are challenging to capture with analytical models.
@@
In our work, we seek alternative methods to determine the noise profile of a QC when its properties are unknown. It is desirable to develop characterization techniques with minimal overheads in terms of quantum processor unit (QPU) runtime. We are inspired by prior art in multi-programming  [1] and ensemble distributions [6] to explore how parallelization during runtime can be employed for 1) learning quantum noise / hardware characterization, 2) fidelity improvements in program outcomes, and 3) reductions in runtime latency. This work presents a novel framework that combines the power of multi-programming with Pauli check sandwiching (PCS) for QC characterization and error mitigation. Multi-programming is a technique for parallelized quantum circuit execution while PCS is an error mitigation strategy that utilizes post-selection to reduce errors in quantum circuit outcomes. Our techniques intelligently embed PCS into quantum applications. These programs are then simultaneously executed across a QPU for noise profile characterization. Resulting characterization data adaptively mitigates noisy outcomes within individual thread outcomes. Knowledge of the QPU’s best regions creates a weighted ensemble result with non-trivial gain, found to be up to a $25\%$ fidelity improvement in our reported results.
Figure 1: Example QPU with 20 regions (10 qubits/region) of unique 1q error rates, $p\in[0.0005,0.01]$, and 2q errors, $2p$. PCS-based characterization with multi-programmed circuits learns the QPU’s unknown noise properties.
II Pauli Check Sandwiching
@@
We rely on a technique called ‘Pauli Check Sandwiching’ (PCS) to detect and mitigate errors [2, 4, 3]. In our work, we also find that PCS is a powerful tool for guiding QC characterization. As seen in Fig. 2, PCS surrounds a payload circuit, $U$, with controlled Pauli operator checks. Errors on $U$ can be detected on an ancilla through phase kickback. It is important that the relationship $R_{1}UL_{1}=U$ holds so the introduced checks do not disturb the original payload circuit semantics.
Recognizing that all errors can be decomposed into Pauli operators is a key concept in understanding PCS. By cleverly constructing Pauli sandwich checks dependent on $U$, anti-commutativity relations between the errors and checks can reveal errors in the ancilla qubits. Measuring 1 in any ancillas indicates that phase kickback occurred, meaning anti-commutation between the Paulis, and thus, a present error. As part of the PCS protocol, we discard this error-corrupted shot, leading to an increase in final distribution fidelity of $U$.
Practical application of PCS requires consideration of tradeoffs. First, determining the Pauli unitary used in the check increases in complexity with the circuit. Second, Pauli check logic must be decomposed into the supported gate set of the QC. Third, the overhead of adding PCS into the circuit must not outweigh the corrective benefit in terms of gate error. We consider these constraints in our solution.
@@
III QC Characterization and Error Mitigation with PCS
With the PCS protocol, we can 1) learn the noise profile of the QC and 2) increase circuit fidelity by eliminating the states with errors. First, we utilize single-qubit checks only on edge qubits to minimize PCS gate count overhead. We then maximally utilize hardware via multi-programming; a QC with $n$ physical qubits holds $threads=\lfloor n/(q_{algorithm}+q_{ancilla})\rfloor$ instances of a circuit running in parallel. The value of the checks is used during post processing to filter errored outcomes from each of the threads, creating vectors of error-mitigated counts, $c_{i}$. The percentage of discarded states, $d_{i}$, gives insight into the underlying noise profile of the QC when they are compared. Each local $d_{i}$ is then used as a weight that scales $c_{i}$ to create $s_{i}=c_{i}*\frac{min(d)}{d_{i}}$. A cumulative distribution, $S$, is created from the sum of the $s_{i}$ vectors from the PCS-protected local threads, $S=\sum s_{i}$.
IV Experimental Results
We simulated the fidelity rates of PCS in both a GHZ / sensing circuit and Toffoli circuit to concretely measure the improvement. We expand the Fig. 1 model QC and study a system with 60 10-qubit regions where single-qubit depolarizing noise $p\in[0.0005,0.03]$ and two-qubit gates are $2p$. We find that PCS not only mitigates error in individual circuit distributions – check data can also be used to intelligently combine the results of multiple threads cross-QPU, demonstrating measurable noise reductions in final outcomes. Fig. 3 shows our results, detailing a multiple percentage point increase in fidelity when multi-programmed applications are protected with PCS and a cumulative distribution is created from local results weighted by check data. Even with only the edge qubits protected, a measurable benefit results. Fig. 4 documents the relationship between discarded shots and error rate, illustrating PCS’s potential for QC error characterization. We believe this suggests that PCS has additional promise to guide efficient qubit mapping.
@@ IV Experimental Results
Figure 3: Fidelity of circuits with no PCS (base) vs. PCS protection. 10,000 shots were used for both circuits in each simulator region (60 total), and fidelity improvement of the PCS weighted ensemble was determined to be $25\%$ (left) and $16.25\%$ (right) as compared to a naive ensemble.
@@
Figure 4: Shots removed / total shots as error rates increase. 10,000 shots on a GHZ circuit were executed for every error rate (60 total).
V Acknowledgments